OpenStudy (anonymous):

1.)In her catering business, Elena usually estimates that 10 pounds of meat can feed about 36 people. Write a proportion that can be used to determine how many pounds of meat should she prepare if she is expecting 162 people, then solve. 2.) A repair service charges $60.00 to come to your home plus $19.50 per hour. If y represents the total charge and n represents the number of hours worked, write an equation that can be used to represent the total charge for a repair job? Then solve if the repair man worked for 4 1/2 hours.

OpenStudy (anonymous):

@xmoses1

OpenStudy (anonymous):

he might be able to help. :)

OpenStudy (littlebird):

1) 10/36=x/162 Solve for x 2) y=19.5n+60 (plug in 4.5 where n is)